8. If the real normalized functions f(x) and g(x) are not orthogonal, show that their sum f(x) +g(x) and their difference f(x)−g(x) are orthogonal.


1
Expert's answer
2021-12-27T13:03:16-0500

Inner product for real functions: "\langle f,g\rangle = \int\limits_{-\infty}^{\infty }f(x)g(x)dx" .


If "f(x)" and "g(x)" are normalized, then "\langle f,f\rangle = \int\limits_{-\infty}^{\infty }f(x)^2dx=1" and "\langle g,g\rangle = \int\limits_{-\infty}^{\infty }g(x)^2dx=1" .


We need to prove, that "\langle f+g, f-g\rangle =0" .


"\langle f+g,f-g\rangle = \int\limits_{-\infty}^{\infty }\big(f(x)+g(x)\big)\cdot \big( f(x)-g(x)\big)dx=

\int\limits_{-\infty}^{\infty }\big(f(x)^2-g(x)^2\big)dx= \int\limits_{-\infty}^{\infty }f(x)^2dx-\int\limits_{-\infty}^{\infty }g(x)^2dx=1-1=0."


So, "f(x)+g(x)" and "f(x)-g(x)" are orthogonal